# Classical origins of Landau-incompatible transitions

### v1

@@ Abstract
The microscopic symmetries in Eq. 2 for both classical and quantum Hamiltonians in Eqs. 1 and 17 are anomaly-free at all scales. This is verified by the presence of the symmetry-allowed operator $\cos(\phi)$ in Eq. 5 which produces a trivial phase — a sufficient condition for the absence of anomalies. However, when $\cos(\phi)$ is irrelevant, a new continuous $\phi$ rotation symmetry emerges which has a well-known mixed anomaly with the microscopic ${\mathcal{R}}$ symmetry [9, 36]. We may say that this microscopic-emergent mixed anomaly endows stability to the exotic transitions. A case was made for a possible transition between the Néel to valence-bond solid (VBS) phases on the honeycomb lattice [37] being stabilized by a similar anomaly in Ref [9]. However, some care is warranted. In fact, the critical Ising model, which is the archetypal Landau-compatible transition, also has a mixed anomaly between the microscopic Ising symmetry and the emergent Kramers-Wannier symmetry [38]. The key difference is that the microscopic-emergent mixed anomaly of the critical Ising model is explicitly broken by relevant operators, whereas the same is broken by irrelevant operators in our model. In summary, although all microscopic symmetries are anomaly-free, a mixed anomaly between microscopic and emergent symmetries that is preserved by all relevant operators can stabilize both Landau-incompatible and unnecessary critical transitions. It is unclear if this is a necessary condition for Landau-incompatible transitions in general.
@@
Outlook: We have investigated a classical 2d statistical mechanical model hosting a stable Landau-incompatible transition and unnecessary criticality. These transitions are driven by the melting of charged defects and stabilized by a mixed anomaly between microscopic and emergent symmetries unbroken by relevant operators despite all microscopic symmetries being anomaly-free.
Our work opens up several lines of future investigation. An obvious one being whether we can find similar phenomena in higher dimensional and/or quantum models. In particular, the existence of the Landau-incompatible Néel to valence-bond-solid transition is known to have a strong dependence on the underlying lattice, and it would be interesting to find alternative settings where such a transition can be present on any lattice, similar to our models. It would be useful to further clarify if the stable mixed microscopic-emergent anomaly is a necessary condition for Landau-incompatible transitions. For odd $Q$ models, it would be interesting to find an explicit deformation that terminates the unnecessary critical lines. In [12], it was conjectured that unnecessary criticality in quantum models is associated with stable edge modes. In our classical model, we expect this to take the form of boundary magnetization. We expect non-trivial boundary phenomena in the vicinity of the Landau-incompatible transitions, also driven by the same mechanism that binds charges to the domain walls. It would be interesting to study them using tools from boundary criticality [39]. Finally, we have shown that exotic transitions can exist under relatively ordinary conditions. Thus it may be simpler than previously imagined to observe this phenomena experimentally, and possible settings should be investigated.
@@
Acknowledgments: The authors thank Ruben Verresen, Ryan Thorngren, Steve Simon, Paul Fendley, Sounak Biswas and Yuchi He for useful discussions. The work of A.P. is supported by the European Research Council under the European Union Horizon 2020 Research and Innovation Programme, Grant Agreement No. 804213-TMCS.
Appendix A Spontaneous symmetry breaking and Landau-compatibility
(a) $G=D_{4}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$
(b) $G=D_{8}\cong{\mathbb{Z}}_{4}\rtimes\mathbb{Z}_{2}$
Figure S1: Hasse diagram of the lattice of conjugate subgroups for the (a) abelian $D_{4}\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and (b) non-abelian $D_{8}\cong{\mathbb{Z}}_{4}\rtimes\mathbb{Z}_{2}$ groups. The arrows represent an inclusion map, and two entries connected by a sequence of maps represent Landau-compatible phases. Horizontal lines represent the various vacua over which their residual symmetries are listed. We see that the phase with all symmetries broken (bottom) and all symmetries preserved (top) are Landau-compatible with all other phases.
A.1 Classifying spontaneous symmetry breaking phases
We begin with a careful discussion of spontaneous symmetry breaking and Landau compatibility. For simplicity, we will assume that we are working with a classical statistical mechanical system whose symmetries $g\in G$ form a group of order $|G|$. In particular, we will not discuss continuous groups nor generalized symmetries [42].
The system is said to be in a spontaneously symmetry broken (SSB) phase if it has multiple vacua that are not invariant under the full set of symmetries. For a given vacuum $v_{\alpha}$, let us denote $H_{\alpha}\subset G$ to be the subgroup of residual symmetries that leaves it invariant. Symmetries in $G$, but not in $H_{\alpha}$, then transform between the vacua.
Using that symmetries in the set $G\setminus H_{\alpha}$ give us transformations between different vacua, it is straightforward to see that the transformations from $v_{\alpha}$ to other vacua $v_{\beta}$ are labelled by cosets of $H$. Hence, we can reach $[G:H_{\alpha}]$ other vacua this way  [^2] In general there may be more—consider the limit of two decoupled systems both exhibiting SSB independently.; and if $g_{\alpha\beta}$ takes $v_{\alpha}\rightarrow v_{\beta}$, we have that the symmetry group of $v_{\beta}$ is $H_{\beta}=g_{\alpha\beta}H_{\alpha}g_{\alpha\beta}^{-1}$. This perspective gives us the family of vacua generated by the vacuum with residual symmetry $H_{\alpha}$, and considering the different cosets will give us the same family independent of the initial choice of $v_{\alpha}$. We conclude that the different possible symmetry-breaking phases of a system with a symmetry group $G$ are classified by all possible subgroups, up to conjugation  [^3] The identification by conjugation is often not stated in literature [41]. However, this identification is important for correct classification, and without it we would overcount the number of SSB phases. For abelian groups, conjugation acts trivially, and various subgroups indeed represent distinct phases. For non-abelian groups, this is not the case..
For a finite group $G$, the different SSB phases are nicely organized by the lattice of conjugate subgroups and visualized by a Hasse diagram where the families of conjugate subgroups are connected by the presence of an inclusion map, i.e., when one family is a subgroup of the other. In Fig. S1, we have shown this for the Dihedral symmetries $D_{2Q}\cong{\mathbb{Z}}_{Q}\rtimes\mathbb{Z}_{2}$ considered in the main text for $Q=2,4$ with the presentation,
$D_{2Q}=\langle{\mathcal{R}},{\mathcal{T}}|{\mathcal{R}}^{Q}={\mathcal{T}}^{2}=1,~{}{\mathcal{R}}{\mathcal{T}}={\mathcal{T}}{\mathcal{R}}^{-1}\rangle.$ (19)
A.2 Landau compatibility and incompatibility
We can also distinguish between Landau-compatible and incompatible transitions using the lattice of conjugate subgroups and its Hasse diagram. Two SSB phases represented by two families of conjugate subgroups are Landau-compatible if they are connected in the Hasse diagram by the composition of a sequence of arrows. Physically, we can understand this as follows: If we place ourselves in one of the vacua of the SSB phase with residual symmetries $H_{\alpha}\subset G$, we can treat it as a standalone system with symmetries $H_{\alpha}$. These can be spontaneously broken into a conjugate family of subgroups $K\subset H_{\alpha}$. All other transitions are Landau incompatible. Physically, a transition between Landau-incompatible phases cannot be understood by a hierarchical splitting of each vacuum, but rather by a more drastic process involving multiple vacua coming together and reorganizing themselves.
In particular the fully symmetric phase with unique vacuum and fully broken phase with $|G|$ vacua are Landau compatible with all SSB phases. The interesting cases are the phases with partial symmetry breaking as we saw in the main text. Landau-compatible transitions are characterized by a change in the number of vacua (although this change does not guarantee compatibility). Moreover, Landau-incompatible transitions can occur between two SSB phases with the same number of vacua, as seen in the main text.
Appendix B More details of the phase diagrams
@@
The important symmetries form the group $D_{2Q}\cong{\mathbb{Z}}_{Q}\rtimes\mathbb{Z}_{2}$ with the presentation in Eq. 19 and the following action on the planar rotors:
@@
The phase diagrams of Eq. 20 are produced by replacing Eqs. 20 and 17 with the effective continuum gaussian theory in either of the two dual forms
@@
and tracking the scaling dimensions of operators $\cos(\phi)$ and $\cos(Q\ell\theta)$. Setting $h=\gamma_{\ell}=0$, these can be determined directly from Eq. 22 to be
$[\cos(\phi)]=\pi\beta,~{}[\cos(\ell Q\theta)]=\frac{\ell^{2}Q^{2}}{4\pi\beta}.$ (23)
However, for $\gamma_{\ell},~{}h$ non-zero, these cannot be determined exactly. Nevertheless, when the theory flows to a gaussian theory, the following relation satisfied by Eq. 23 and stated in the main text holds
$[\cos(\phi)][\cos(\ell Q\theta)]=\ell^{2}Q^{2}/4.$ (24)
In this case, we can parametrize the theory using a single parameter which we take to be the scaling dimension of $[\cos(\phi)]=K$, that determines all other scaling dimensions as well as correlation functions
$[\cos(\ell Q\theta)]=\frac{\ell^{2}Q^{2}}{4K},~{}\langle\cos(\phi(x))\cos(\phi(y))\rangle\sim|x-y|^{-2K},~{}\langle\cos(\ell Q\theta(x))\cos(\ell Q\theta(y))\rangle\sim|x-y|^{-\frac{\ell^{2}Q^{2}}{2K}}.$ (25)
$K$ is a function of all parameters $\{\gamma_{\ell}\}$ and $T$.
The various phases present in the model are:
@@
    A trivial paramagnet driven by $\cos(\phi)$.
@@
    Partial symmetry-breaking regions (p-SSB±) driven by $\gamma_{1}\cos(Q\theta)$ for $\gamma_{1}>0$ and $\gamma_{1}<0$ with vacua
@@
    Full-Symmetry Breaking Phase (f-SSB) driven by $\gamma_{2}\cos(2Q\theta)$ for $\gamma_{2}<0$ with vacua
@@
  • $T^{\alpha}_{Q}$, when $\cos(2Q\theta)$ is marginal, i.e., $K=Q^{2}/2$,
  • $T^{\beta}_{Q}$ when $\cos(Q\theta)$ is marginal, i.e., $K=Q^{2}/8$ and
  • $T^{\gamma}$ when $\cos(\phi)$ is marginal, i.e., $K=2$.
We may ask if there are other perturbations missed in the models we have considered that can qualitatively change the phase diagram. It is easy to verify that there are no other primaries [19] that are allowed by symmetry. For example, $\sin(Q\theta)$, which could connect the $\gamma_{1}<0$ and $\gamma_{1}>0$ p-SSB± regions, is disallowed by the $\theta\mapsto-\theta$ symmetry. But what about descendants [19]? In particular for $Q\geq 3$, for $T^{\alpha}<T<\min(T^{\beta}_{Q},T^{\gamma})$ the only relevant symmetry-allowed primary operator in Eq. 22 is $\cos(Q\theta)$. Discounting the symmetry-disallowed operators, we should consider $\partial_{x}\left(\cos(Q\theta)\right)$ with dimension $[\cos(Q\theta)]+1$. This is symmetry-allowed and would be relevant when $[\cos(Q\theta)]<1$. However, this operator is a boundary term and would not affect the bulk phase diagram. We do not explore the interesting question of how the presence of descendants can affect boundary phenomena in this work.
@@
for $|\Delta|<1$ by identifying $K=2\arccos\Delta$ via Bethe ansatz [45].
B.1 Double frequency sine-Gordon and f-SSB to p-SSB± Ising transition
@@
Figure S2: The Ising transition for $\gamma_{1}>0$ seen from the potential  Eq. 30 in the double-frequency sine-Gordon model in Eq. 29. We see a transition at $\gamma_{1}=4|\gamma_{2}|$
@@
For $\gamma_{2}<0$, $V(\theta)$ has a unique minimum for large $|\gamma_{1}|$ at $\theta=0$ (for $\gamma_{1}>0$) and $\theta=\pi$ (for $\gamma_{1}<0$). As we reduce $|\gamma_{1}|$, we see in Fig. S2 that the $\theta\mapsto-\theta$ symmetry is spontaneously broken in the vacuum at $\gamma_{1}=4|\gamma_{2}|$. Near this point, we can Taylor expand Eq. 29 to get (after appropriate redefinitions) a 2d real scalar field theory with Ising symmetry
@@
B.2 Self-dual sine-Gordon model and PM to p-SSB± transition
@@
$S=\int d^{2}x\left[\frac{\left(\nabla\phi\right)^{2}}{4\pi Q}-g\left(\cos(\phi)+\cos Q\theta\right)\right].$ (32)
For $Q\leq 4$, when $\cos(\phi)$ and $\cos Q\theta$ are both not irrelevant, the fate of Eq. 32 under RG flow was analyzed in Ref. [13]. Their results are summarized as follows:
@@
    For $Q=4$, $\cos(\phi)$ and $\cos Q\theta$ are both marginal and induce a flow in the $c=1$ conformal manifold [46], along the orbifold branch which describes the scaling limit of the Ashkin-Teller model [47]. Thus, the transition has varying critical exponents.
The RG flow of Eq. 32 is trivial for $Q\geq 5$ when $\cos(\phi)$ and $\cos Q\theta$ are irrelevant, giving us a $c=1$ gaussian theory. This represents an intermediate gapless phase rather than a direct transition between PM and p-SSB± phases.
B.3 Phase diagrams for $\gamma_{2}>0$
@@ B.4 $Q=1$ phase diagrams
Figure S3: Phase diagrams for the $Q=1$ Hamiltonian of Eq. 20. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines.
Let us begin with the trivial, yet instructive case of the phase diagram for $Q=1$. The only non-trivial symmetry is ${\mathcal{T}}$ in Eq. 21 and generates a $\mathbb{Z}_{2}$ group.
@@
An interesting observation is the absence of a direct phase transition between the $\cos\theta$ and $\cos(\phi)$ dominant regions. This was argued from the analysis of the RG flow of the self-dual sine-Gordon model in Eq. 32 with $Q=1$. A much simpler way to see the same fact is to consider the $Q=1$ quantum Hamiltonian in Eq. 28
@@
By setting $\gamma_{1}\rightarrow\infty$, we get the $\cos\theta$ dominant phase with a product ground state. This is clearly adiabatically connected to the ground state of $h\sum_{j}(-1)^{j}S^{z}_{j}$ which is equivalent to the $\cos(\phi)$ dominated high temperature phase of the classical model.
We see that the nature of the phase diagram depends on the sign of $\gamma_{2}$, as discussed before and shown in Fig. S3(a,c). For $\gamma_{2}=0$, the phase diagram is modified as shown in Fig. S3(b). The Ising and first order lines can be analyzed using the double-frequency sine-Gordon potential
@@
and repeating the analysis in Section B.1. The 3-state Potts universality class resulting where the two lines meet at $T^{\beta}_{3}$ and $\gamma_{1}=0$ obtains from the flow of the self-dual sine-Gordon model in Eq. 32 with $Q=3$.
If we further tune $\gamma_{3}=0$ (not shown in Fig. S3), we get a continuous unnecessary critical line in the phase diagram for $T<T^{\gamma}$ along $\gamma_{1}=0$. This is not stable and needs fine-tuning three parameters, unlike for $Q\geq 3$ where the unnecessary criticality needs fine-tuning only one parameter, i.e, $\gamma_{1}=0$.
B.5 $Q=2$ phase diagrams

### v2

@@ Abstract
All microscopic symmetries in Eq. 2 are anomaly-free. This is verified by the presence of the symmetry-allowed operator $\cos(\phi)$ in Eq. 4 which produces a trivial phase—a sufficient condition for the absence of anomalies. What we have in our model is arguably a more pedestrian route for the binding of charges to defects—such a bound state may find itself energetically more favourable  [^1] An analogous charge-binding transition and associated exotic critical phenomena in a 2d classical model were investigated in Refs [70, 71]. However, when $\cos(\phi)$ is irrelevant, a new continuous $\phi$ rotation symmetry emerges that is preserved by all remaining relevant operators  [^2] The critical Ising model, which is the archetypal Landau-compatible transition, also has a mixed anomaly between the microscopic Ising symmetry and the emergent Kramers-Wannier symmetry [72]. The key difference is that this mixed anomaly is explicitly broken by relevant operators, whereas the same is broken by irrelevant operators in our model. which has a mixed anomaly with the microscopic ${\mathcal{R}}$ symmetry [12, 46]. This microscopic-emergent mixed anomaly may be said to stabilize the Landau-incompatible and unnecessary critical transitions. It is unclear if this is a necessary precondition.
@@
Outlook: We have investigated classical 2d statistical mechanical models hosting stable Landau-incompatible transitions and unnecessary criticality. These transitions are driven by the melting of charged defects and stabilized by a mixed anomaly between microscopic and emergent symmetries unbroken by relevant operators despite all microscopic symmetries being anomaly-free.
Our work opens several lines of future investigation. An obvious one is whether we can find other classical models that exhibit similar phenomena  [^3] Numerical evidence for a direct Landau-incompatible transition in an alternative classical model was presented in Ref. [73], especially in higher dimensions. The archetype DQC transition, between Néel to valence-bond-solid phases [2] has recently been shown to be first-order in nature [50] and it would be interesting to find alternative, classical settings where a Landau-incompatible transition can be present between other phases. For odd $Q$ models, it would be interesting to see how the unnecessary critical surface terminates. In [15, 51], it was argued that unnecessary criticality in quantum models is stabilized by the encircling states forming a non-trivial family  [52, 53, 54]. We expect this to be true for our classical model and it would be most interesting to explore this connection further. It would also be useful to further clarify if the stable mixed microscopic-emergent anomaly is a necessary condition for the exotic transitions studied here.
We have shown that exotic transitions can exist under relatively ordinary conditions within reach of existing experiments. It would be gratifying to validate this in more experimental setups. Finally, it would be illuminating to explore what other phenomena attributed to quantum fluctuations can have a classical origin.
@@
Acknowledgments: The authors thank Ruben Verresen, Ryan Thorngren, Steve Simon, Paul Fendley, Sounak Biswas, Yuchi He and Saranesh Prembabu for useful discussions and the anonymous referees for insightful comments. A.P. was supported by the European Research Council under the European Union Horizon 2020 Research and Innovation Programme, Grant Agreement No. 804213-TMCS.
Appendix I More details of the phase diagrams
@@
The important symmetries form the group $D_{2Q}\cong{\mathbb{Z}}_{Q}\rtimes\mathbb{Z}_{2}$ and have the following action on the planar rotors:
@@
The phase diagrams for each $Q$ are produced by replacing Eq. 19 with the effective continuum gaussian theory in either of the two dual forms
@@
and tracking the scaling dimensions of operators $\cos(\phi)$ and $\cos(Q\ell\theta)$. Let us begin by setting all $\gamma_{\ell}=0$, giving the familiar isotropic XY model in Eq. 19, whose phase diagram is reproduced by the gaussian model perturbed by a single operator $\cos(\phi)$ in Eq. 21. It is well known that there exists a critical temperature $T_{\gamma}$ such that for $T>T_{\gamma}$, $\cos(\phi)$ is relevant and drives the system to the trivial, disordered phase and for $T<T_{\gamma}$, it is irrelevant and Eq. 21 flows to the critical phase with pseudo long-range order described by the following effective theory
$S\rightarrow\int d^{2}x\frac{\left(\nabla\phi\right)^{2}}{8\pi^{2}K_{\rm eff}}\leftrightarrow\int d^{2}x\frac{K_{\rm eff}}{2}\left(\nabla\theta\right)^{2}.$ (22)
In this case, we can parametrize the theory using a single parameter $K_{\rm eff}$ which is related to the scaling dimension of $\cos(\phi)$. This determines all other scaling dimensions as well as correlation functions [13]
$[\cos(\phi)]=\pi K_{\rm eff},~{}[\cos(\ell Q\theta)]=\frac{\ell^{2}Q^{2}}{4\pi K_{\rm eff}},$ (23)
$\langle\cos(\phi(x))\cos(\phi(y))\rangle\sim|x-y|^{-2\pi K_{\rm eff}},~{}\langle\cos(\ell Q\theta(x))\cos(\ell Q\theta(y))\rangle\sim|x-y|^{-\frac{\ell^{2}Q^{2}}{2\pi K_{\rm eff}}}.$ (24)
Let us now introduce $\gamma_{\ell}\neq 0$. For small values, their renormalization group (RG) flow equations are also determined from $K_{\rm eff}$ as follows
$\frac{d\gamma_{\ell}}{ds}\approx\left(2-\frac{\ell^{2}Q^{2}}{4\pi K_{\rm eff}}\right)\gamma_{\ell},$ (25)
where $s$ is the length-scale of the RG flow. We see that when $\frac{\ell^{2}Q^{2}}{4\pi K_{\rm eff}}<2$, $\gamma_{\ell}$ is relevant and grows at large distances whereas for $\frac{\ell^{2}Q^{2}}{4\pi K_{\rm eff}}>2$, $\gamma_{\ell}$ is irrelevant and shrinks. When multiple $\gamma_{\ell}$ are relevant, so long as they are all of the same order, the operators with the smallest scaling dimensions dominate (in this case corresponding to smallest $\ell$). Since we tune $\gamma_{1}$ keeping all other $\gamma_{\ell}$ fixed, we only need to keep track of $\ell=1,2$ to determine our phase diagram of interest. The various phases present in the model and their regimes are:
@@
    A trivial disordered paramagnet driven by $\cos(\phi)$.
@@
    Partial symmetry-breaking regions (p-SSB±) when $\cos(Q\theta)$ is relevant, driven by $\gamma_{1}\cos(Q\theta)$ for $\gamma_{1}>0$ and $\gamma_{1}<0$ with vacua
@@
    Full symmetry-breaking phase (f-SSB) when $\cos(2Q\theta)$ is relevant, driven by $\gamma_{2}\cos(2Q\theta)$ for $\gamma_{2}<0$ with vacua
@@
  • $T^{\alpha}_{Q}$, when $\cos(2Q\theta)$ is marginal, i.e., $K_{\rm eff}=Q^{2}/(2\pi)$,
  • $T^{\beta}_{Q}$ when $\cos(Q\theta)$ is marginal, i.e., $K_{\rm eff}=Q^{2}/(8\pi)$ and
  • $T^{\gamma}$ when $\cos(\phi)$ is marginal, i.e., $K_{\rm eff}=2/\pi$.
For $Q\geq 3$, along the $\gamma_{1}=0$ line, for the range of temperatures between $T_{\alpha}^{Q}$ and $T^{\beta}_{Q}$, we get the Landau-incompatible transitions (even $Q$) and unnecessary critical line (odd $Q$) when $\cos(Q\theta)$ is the only relevant operator.
We may ask if there are other perturbations missed in the models we have considered that can qualitatively change the phase diagram. It is easy to verify that there are no other primary [22] scaling operators that are allowed by symmetry. For example, $\sin(Q\theta)$, which could connect the $\gamma_{1}<0$ and $\gamma_{1}>0$ p-SSB± regions, is disallowed by the $\theta\mapsto-\theta$ symmetry. But what about descendant operators [22]? In particular for $Q\geq 3$, for $T^{\alpha}<T<\min(T^{\beta}_{Q},T^{\gamma})$ the only relevant symmetry-allowed primary operator in Eq. 21 is $\cos(Q\theta)$. Discounting the symmetry-disallowed operators, we should consider $\partial_{x}\left(\cos(Q\theta)\right)$ with dimension $[\cos(Q\theta)]+1$. This is symmetry-allowed and would be relevant when $[\cos(Q\theta)]<1$. However, this operator is a boundary term and would not affect the bulk phase diagram. In this work, we do not explore the interesting question of how the presence of descendants can affect boundary phenomena.
@@
for $|\Delta|<1$ by identifying $K_{\rm eff}=2\arccos\Delta$ via Bethe ansatz [55].
I.1 Double frequency sine-Gordon and f-SSB to p-SSB± Ising transition
@@
Figure 7: The Ising transition for $\gamma_{1}>0$ seen from the potential  Eq. 30 in the double-frequency sine-Gordon model in Eq. 29. We see a transition at $\gamma_{1}=4|\gamma_{2}|$
@@
For $\gamma_{2}<0$, $V(\theta)$ has a unique minimum for large $|\gamma_{1}|$ at $\theta=0$ (for $\gamma_{1}>0$) and $\theta=\pi$ (for $\gamma_{1}<0$). As we reduce $|\gamma_{1}|$, we see in Fig. 7 that the $\theta\mapsto-\theta$ symmetry is spontaneously broken in the vacuum at $\gamma_{1}=4|\gamma_{2}|$. Near this point, we can Taylor expand Eq. 29 to get (after appropriate redefinitions) a 2d real scalar field theory with Ising symmetry
@@
I.2 Self-dual sine-Gordon model and PM to p-SSB± transition
@@
$S=\int d^{2}x\left[\frac{\left(\nabla\phi\right)^{2}}{4\pi Q}-g\left(\cos(\phi)+\cos(Q\theta)\right)\right].$ (32)
For $Q\leq 4$, when $\cos(\phi)$ and $\cos(Q\theta)$ are both not irrelevant, the fate of Eq. 32 under RG flow was analyzed in Ref. [17]. Their results are summarized as follows:
@@
    For $Q=4$, $\cos(\phi)$ and $\cos(Q\theta)$ are both marginal and induce a flow in the $c=1$ conformal manifold [56], along the orbifold branch which describes the scaling limit of the Ashkin-Teller model [57]. Thus, the transition has varying critical exponents.
The RG flow of Eq. 32 is trivial for $Q\geq 5$ when $\cos(\phi)$ and $\cos(Q\theta)$ are irrelevant, giving us a $c=1$ gaussian theory. This represents an intermediate gapless phase rather than a direct transition between PM and p-SSB± phases.
I.3 Phase diagrams for $\gamma_{2}>0$
@@ I.4 $Q=1$ phase diagrams
Figure 8: Phase diagrams for the $Q=1$ Hamiltonian of Eq. 19. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines.
Let us begin with the trivial, yet instructive case of the phase diagram for $Q=1$. The only non-trivial symmetry is ${\mathcal{T}}$ in Eq. 20 and generates a $\mathbb{Z}_{2}$ group
@@
An interesting observation is the absence of a direct phase transition between the $\cos(\theta)$ and $\cos(\phi)$ dominant regions. This was argued from the analysis of the RG flow of the self-dual sine-Gordon model in Eq. 32 with $Q=1$. A much simpler way to see the same fact is to consider the $Q=1$ quantum Hamiltonian in Eq. 28
@@
By setting $\gamma_{1}\rightarrow\infty$, we get the $\cos(\theta)$ dominant phase with a product ground state. This is clearly adiabatically connected to the ground state of $h\sum_{j}(-1)^{j}S^{z}_{j}$ which is equivalent to the $\cos(\phi)$ dominated high temperature phase of the classical model.
We see that the nature of the phase diagram depends on the sign of $\gamma_{2}$, as discussed before and shown in Fig. 8(a,c). For $\gamma_{2}=0$, the phase diagram is modified as shown in Fig. 8(b). The Ising and first order lines can be analyzed using the double-frequency sine-Gordon potential
@@
and repeating the analysis in Section I.1. The 3-state Potts universality class where the two lines meet at $T^{\beta}_{3}$ and $\gamma_{1}=0$ is obtained from the flow of the self-dual sine-Gordon model in Eq. 32 with $Q=3$.
If we further tune $\gamma_{3}=0$ (not shown in Fig. 8), we get a continuous unnecessary critical line in the phase diagram for $T<T^{\gamma}$ along $\gamma_{1}=0$. This is not stable and needs fine-tuning three parameters, unlike for $Q\geq 3$ where the unnecessary criticality needs fine-tuning only one parameter, i.e, $\gamma_{1}=0$.
I.5 $Q=2$ phase diagrams